OpenStudy (anonymous):

what is the derivative of 3x^2/3?

OpenStudy (hunus):

The derivative of equations of the form \[f(x)=c*x^{n}\] where c is a constont are equal to \[f'(x)=c*nx^{n-1}\] Try that on your equation

OpenStudy (anonymous):

so i would do 3*2/3*x^2/3-1?

OpenStudy (hunus):

Yes

OpenStudy (anonymous):

ok so 2x^1/3

OpenStudy (hunus):

Close. 2/3 - 1 = -1/3

OpenStudy (anonymous):

if i want to solve for x. and the equation is 2x^-1/3=0 how do i get rid of the -1/3

OpenStudy (hunus):

You don't really need to. Because x=0 will give you a zero no matter what power x is raised to

OpenStudy (hunus):

You could raise both powers to the 1/3 power \[(2x^{-1/3})^{1/3} = 0^{1/3}\] \[(2x^{-1/3})^{-3} = 0^{-3}\] \[2x = 0\]